# Sharp Analysis for KL-Regularized Contextual Bandits and RLHF

### v1

@@ Appendix C Proofs from Section 3
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{\widehat{\theta}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\widehat{\theta}}^{\eta}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\ln Z_{\theta_{*}}^{\eta}(x)\big{]}-\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\ln Z_{\widehat{\theta}}^{\eta}(x)\big{]}-\mathbb{E}_{x\sim d_{0}}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\big{(}R(\theta_{*},x,a)-R(\widehat{\theta},x,a)\big{)}\bigg{]}$
For an arbitrary reward function $f:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, let $\Delta(x,a)=f(x,a)-R(\theta_{*},x,a)$. Consider the following first derivative of $J(f)=\ln Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta(x,a)$, where $Z_{f}^{\eta}(x)=\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$ and $\pi_{f}^{\eta}(a|x)\propto\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$.
$\quad\frac{\partial}{\partial\Delta(x,a)}\bigg{[}\ln Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta(x,a)\bigg{]}$
@@
Therefore, there exists $f(\cdot,\cdot)=\gamma R(\widehat{\theta},\cdot,\cdot)+(1-\gamma)R(\theta_{*},\cdot,\cdot)$ such that  ($\gamma\in(0,1)$)
@@
From Lemma C.4, if $m\geq 128\eta^{2}D^{2}B^{2}\cdot\log(2N_{\mathcal{R}}(\epsilon_{c})/\delta)$, for any $(x,a)\in\mathcal{X}\times\mathcal{A}$ such that $\pi_{0}(a|x)>0$, it holds that
@@
Let $\epsilon_{c}=\min\{\frac{\epsilon}{(1+c_{m,n}^{-1})B},\frac{1}{8(1+c_{m,n})B\eta^{2}D^{2}},B\}$. From Lemma C.3, if $m\geq 128\eta^{2}D^{2}B^{2}\cdot\log(2N_{\mathcal{R}}(\epsilon_{c})/\delta)$ and $n\geq\eta/\epsilon\cdot B^{2}\log(N_{\mathcal{R}}(\epsilon_{c})/\delta))$ and $n=c_{m,n}m$ then with high probability the output policy $\pi_{\widehat{\theta}}^{\eta}$ is $O(\epsilon)$ optimal. ∎
@@
The proof follows the same lines as Theorem 4.4 by replacing the global-coverage condition with the local-coverage condition. It still holds that
@@
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})\leq\eta\cdot C_{\rho_{\mathrm{KL}}}\cdot O\big{(}\frac{1}{n}B\log(N_{\mathcal{R}}(\epsilon_{c})/\delta)+B(1+c_{m,n}^{-1})\epsilon_{c}\big{)}$
@@ Appendix D Proofs from Section 4
The proof follows a similar construction as the one for Theorem 3.1. Consider a simple case when $|\mathcal{X}|=M$ and $|\mathcal{A}|=2$. We suppose that the context $x$ is drawn uniformly from $\mathcal{X}$ at the beginning of each round. Let $\Theta$ be the set consisting of mappings from $\mathcal{X}$ to $\mathcal{A}=\{0,1\}$. For each $\theta\in\Theta$, we have $R(\theta,x,a)=\begin{cases}c&\text{if }a=\theta(x),\\ 0&\text{if }a\neq\theta(x),\end{cases}$ where $c>0$ is a constant, and $\theta(x)$ is the optimal action under context $x$ when the model is $\theta$.
@@
$|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\leq\sqrt{\log(1/2+e^{c}/4+e^{-c}/4)\mathbb{E}_{\theta_{1}}[N(x_{0})]}\leq\sqrt{c^{2}\mathbb{E}_{\theta_{1}}[N(x_{0})]}=\sqrt{c^{2}T/M},$
@@
Set $A$ to be the event that $\pi_{out}(\theta_{1}(x_{0})|x_{0})>1/2$. Then we have
$\mathbb{P}_{\theta_{1}}(\pi_{out}(\theta_{1}(x_{0})|x_{0})\leq 1/2)+\mathbb{P}_{\theta_{2}}(\pi_{out}(\theta_{2}(x_{0})|x_{0})\leq 1/2)\geq 1-|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\geq 1-\sqrt{c^{2}T/M}.$
@@
$\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{P}_{\theta}(\pi_{out}(\theta(x_{0}))\leq 1/2)\geq\frac{1}{2}-\sqrt{c^{2}T/4M}$
@@
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{out}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\ln\frac{\pi_{out}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{out}}\bigg{[}\ln\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{out}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]},$
where the last equality follows from the fact that $\pi_{\theta_{*}}^{\eta}\propto\pi_{0}(a|x)\cdot\exp(\eta R(\theta_{*},x,a))$.
To bound the suboptimality gap, we further have
$\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$=\mathbb{E}_{\theta\sim Unif(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{E}_{a\sim\pi_{out}(\cdot|x)}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$\geq\mathbb{E}_{\theta\sim Unif(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{P}_{\theta}(\pi_{out}(\theta(x))\leq 1/2)\cdot\biggl{[}\frac{1}{2}\ln\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\ln\frac{1+\exp(\eta c)}{2}\biggr{]}$
$\geq\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/4M}\Bigr{)}\Bigl{[}\frac{1}{2}\ln\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\ln\frac{1+\exp(\eta c)}{2}\Bigr{]}$ (D.1)
@@
$\frac{\text{d}}{\text{d}u}\Bigl{[}\frac{1}{2}\ln\frac{1+e^{-u}}{2}+\frac{1}{2}\ln\frac{1+e^{u}}{2}\Bigr{]}\Big{|}_{u=0}$ $=\frac{1}{2}\Bigl{[}\frac{1}{1+\exp(-u)}-\frac{1}{1+\exp(u)}\Bigr{]}\Big{|}_{u=0}=0,$
$\frac{\text{d}^{2}}{\text{d}u^{2}}\Bigl{[}\frac{1}{2}\ln\frac{1+e^{-u}}{2}+\frac{1}{2}\ln\frac{1+e^{u}}{2}\Bigr{]}$ $=\frac{\exp(u)}{[1+\exp(u)]^{2}}.$
@@
$\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{2}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/4M}\Bigr{)}\eta^{2}c^{2}\cdot\frac{1}{3+\exp(\eta c)}$
@@
$\frac{1}{2}\ln\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\ln\frac{1+\exp(\eta c)}{2}$ $=\frac{1}{2}\ln\frac{2+\exp(\eta c)+\exp(-\eta c)}{4}$
$\geq\frac{1}{2}\cdot\frac{1}{2}\Bigl{(}\ln\frac{\exp(\eta c)+\exp(-\eta c)}{2}\Bigr{)}$
$\geq\frac{1}{4}(\eta c-\ln 2),$ (D.2)
where the first inequality follows from Jensen’s inequality.
Substituting (D.2) into (D.1), we have
$\epsilon\geq\frac{1}{\eta}\mathbb{E}_{\theta\sim Unif(\Theta)}\mathbb{E}_{\pi_{out}}\Bigl{[}\ln\frac{\pi_{out}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{4}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/4M}\Bigr{)}(\eta c-\ln 2)\cdot\frac{1}{\eta}.$
@@
where the equality follows from the fact that $\sigma(r)+\sigma(-r)=1$ and the last inequality holds since $\sigma^{\prime}(r)=\sigma(r)\cdot(1-\sigma(r))\geq e^{-B}$ for all $r\in[-B,B]$. ∎
@@
By Lemma D.1 and Lemma D.2, we have with probability at least $1-\delta$, for any $\theta\in\Theta$,
@@
$0\leq\log(N_{\mathcal{R}}(\epsilon_{c})/\delta)-\frac{1}{8}e^{-B}\sum_{i=1}^{n}\big{(}[R(\theta,x_{i},a_{i}^{2})-R(\theta,x_{i},a_{i}^{1})]-[R(\theta_{*},x_{i},a_{i}^{2})-R(\theta_{*},x_{i},a_{i}^{1})]\big{)}^{2}+O(n\epsilon_{c}).$
@@
for all $\theta^{c}\in\Theta^{c}$.
From Lemma E.2 and the definition of $\Theta^{c}$, we further have
@@
with probability at least $1-2\delta$ from Lemma D.3 and the union bound.
@@
for all $\theta^{c}\in\Theta^{c}$.
From Lemma E.2 and the definition of $\Theta^{c}$, we further have
@@
$\leq O(\frac{1}{m}B^{2}\log(N_{\mathcal{R}}(\epsilon_{c})/\delta)+\frac{1}{m}\sum_{i=1}^{n}\big{(}[R(\widehat{\theta},\widetilde{x}_{i},\widetilde{a}_{i}^{2})-R(\widehat{\theta},\widetilde{x}_{i},\widetilde{a}_{i}^{1})]-[R(\theta_{*},\widetilde{x}_{i},\widetilde{a}_{i}^{2})-R(\theta_{*},\widetilde{x}_{i},\widetilde{a}_{i}^{1})]\big{)}^{2}+B\epsilon_{c}).$ (D.5)
@@
From Lemma D.4, we have with probability at least $1-2\delta$, there exists a mapping $b_{1}:\mathcal{X}\to\mathbb{R}$ such that

### v2

@@ Appendix C Proofs from Section 3
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\widehat{\theta}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\widehat{\theta}}^{\eta}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\log Z_{\theta_{*}}^{\eta}(x)\big{]}-\frac{1}{\eta}\mathbb{E}_{x\sim d_{0}}\big{[}\log Z_{\widehat{\theta}}^{\eta}(x)\big{]}-\mathbb{E}_{x\sim d_{0}}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{\widehat{\theta}}^{\eta}(a|x)\cdot\big{(}R(\theta_{*},x,a)-R(\widehat{\theta},x,a)\big{)}\bigg{]},$
where the first equality follows from the definition of the KL-divergence, the second equality follows from Lemma E.5.
For an arbitrary reward function $f:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, let $\Delta_{f}(x,a)=f(x,a)-R(\theta_{*},x,a)$. Consider the following first derivative of $J(f)=\log Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta_{f}(x,a)$, where $Z_{f}^{\eta}(x)=\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$ and $\pi_{f}^{\eta}(a|x)\propto\pi_{0}(a|x)\cdot\exp(\eta\cdot f(x,a))$.
$\frac{\partial}{\partial\Delta_{f}(x,a)}\bigg{[}\log Z_{f}^{\eta}(x)-\eta\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\cdot\Delta_{f}(x,a)\bigg{]}$
@@
$\quad-\eta\cdot\Delta_{f}(x,a)\cdot\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta\cdot f(x,a)\bigr{)}}{Z_{f}^{\eta}(x)}\cdot\eta+\eta\cdot\Delta_{f}(x,a)\cdot\frac{\bigl{[}\pi_{0}(a|x)\cdot\exp\bigl{(}\eta\cdot f(x,a)\bigr{)}\bigr{]}^{2}}{[Z_{f}^{\eta}(x)]^{2}}\cdot\eta$
$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi_{0}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot f(x,a^{\prime})\bigr{)}}{Z_{f}^{\eta}(x)}\cdot\eta\cdot\Delta_{f}(x,a^{\prime})\cdot\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta\cdot f(x,a)\bigr{)}}{Z_{f}^{\eta}(x)}$
$=-\eta^{2}\pi_{f}^{\eta}(a|x)\Delta_{f}(x,a)+\eta^{2}[\pi_{f}^{\eta}(a|x)]^{2}\cdot\Delta_{f}(x,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{f}^{\eta}(a^{\prime}|x)\pi_{f}^{\eta}(a|x)\Delta_{f}(x,a^{\prime})$
where the first equality is derived by taking the derivative of $\log Z_{f}^{\eta}(x)$ and the second term with respect to $\Delta_{f}$. Therefore, by the Mean Value Theorem, there exists an $f(\cdot,\cdot)=\gamma R(\widehat{\theta},\cdot,\cdot)+(1-\gamma)R(\theta_{*},\cdot,\cdot)$ for some $\gamma\in[0,1]$ such that
@@
where the last inequality holds since
$\sum_{a_{1}\in\mathcal{A}}\sum_{a_{2}\in\mathcal{A}}\pi_{f}^{\eta}(a_{1}|x)\pi_{f}^{\eta}(a_{2}|x)\bigl{(}R(\widehat{\theta},x,a_{1})-R(\theta_{*},x,a_{1})\bigr{)}\bigl{(}R(\widehat{\theta},x,a_{2})-R(\theta_{*},x,a_{2})\bigr{)}$
$=\big{[}\mathbb{E}_{a\sim\pi_{f}^{\eta}(\cdot|x)}[R(\widehat{\theta},x,a)-R(\theta_{*},x,a)]\big{]}^{2}\geq 0.$
Now, we are ready to prove the theorem.
Proof of Theorem 3.8.
By Lemma 3.9, we have
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})$ $\leq\eta\mathbb{E}_{x\sim d_{0}}\Big{[}\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\big{(}R(\widehat{\theta},x,a)-R(\theta_{*}x,a)\big{)}^{2}\Big{]}.$
By Lemma C.4, if $m\geq 128\eta^{2}D^{2}B^{2}\cdot\log(2N_{\mathcal{R}}(\epsilon_{c})/\delta)$, for any $(x,a)\in\mathcal{X}\times\mathcal{A}$ such that $\pi_{0}(a|x)>0$, it holds that
@@
Let $\epsilon_{c}=\min\{\frac{\epsilon}{(1+c_{m,n}^{-1})B},\frac{1}{8(1+c_{m,n})B\eta^{2}D^{2}},B\}$. By Lemma C.3, if $m\geq 128\eta^{2}D^{2}B^{2}\cdot\log(2N_{\mathcal{R}}(\epsilon_{c})/\delta)$ and $n\geq\eta/\epsilon\cdot B^{2}\log(N_{\mathcal{R}}(\epsilon_{c})/\delta)$ and $n=c_{m,n}m$ then with high probability the output policy $\pi_{\widehat{\theta}}^{\eta}$ is $O(\epsilon)$ optimal. ∎
@@
The proof follows the same lines as Theorem 4.7 by replacing the data coverage condition with the local-coverage condition. It still holds that
@@
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})\leq\eta\cdot C_{\rho_{\mathrm{KL}}}\cdot O\bigg{(}\frac{1}{n}B\log(N_{\mathcal{R}}(\epsilon_{c})/\delta)+B(1+c_{m,n}^{-1})\epsilon_{c}\bigg{)}$
@@ Appendix D Proofs from Section 4
The proof follows a similar construction as the one for Theorem 3.6. Consider a simple case when $|\mathcal{X}|=M$ and $|\mathcal{A}|=2$. We suppose that the context $x$ is drawn uniformly from $\mathcal{X}$ at the beginning of each round. Let $\Theta$ be the set consisting of mappings from $\mathcal{X}$ to $\mathcal{A}=\{0,1\}$. For each $\theta\in\Theta$, we have $R(\theta,x,a)=\begin{cases}c&\text{if }a=\theta(x),\\ 0&\text{if }a\neq\theta(x),\end{cases}$ where $c\in(0,1/4)$ is a constant, and $\theta(x)$ is the optimal action under context $x$ when the model is $\theta$.
@@
We have the following upper bound for two Bernoulli distribution $y_{1}\sim Bernoulli(\sigma(c))$ and $y_{2}\sim Bernoulli(\sigma(-c))$ with $\sigma(x)=1/(1+\exp(-x))$:
$\sigma(c)\log\frac{\sigma(c)}{\sigma(-c)}+\sigma(-c)\log\frac{\sigma(-c)}{\sigma(c)}$ $=2\cdot(\frac{1}{2}-\sigma(-c))\log\frac{\sigma(c)}{\sigma(-c)}$
$=\frac{1-e^{-c}}{1+e^{-c}}\cdot\log\frac{1+e^{c}}{1+e^{-c}}$
$\leq\frac{1-e^{-c}}{1+e^{-c}}\cdot\frac{e^{c}-e^{-c}}{1+e^{-c}}$
$\leq\frac{c\cdot(c+e^{\frac{1}{4}}c)}{[(1+e^{-\frac{1}{4}})]^{2}}\leq c^{2},$
where the first equality follows from the fact that $\sigma(-c)=1-\sigma(c)$, the first inequality holds since $\log x\leq x-1$, and the second inequality holds since $c\leq 1/4$.
@@
$|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\leq\sqrt{c^{2}\mathbb{E}_{\theta_{1}}[N(x_{0})]}=\sqrt{c^{2}T/M},$
@@
Set $A$ to be the event that $\pi_{\text{out}}(\theta_{1}(x_{0})|x_{0})>1/2$. Then we have
$\mathbb{P}_{\theta_{1}}(\pi_{\text{out}}(\theta_{1}(x_{0})|x_{0})\leq 1/2)+\mathbb{P}_{\theta_{2}}(\pi_{\text{out}}(\theta_{2}(x_{0})|x_{0})\leq 1/2)\geq 1-|\mathbb{P}_{\theta_{1}}(A)-\mathbb{P}_{\theta_{2}}(A)|\geq 1-\sqrt{c^{2}T/M}.$
@@
$\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{P}_{\theta}(\pi_{\text{out}}(\theta(x_{0}))\leq 1/2)\geq\frac{1}{2}-\sqrt{c^{2}T/4M}$
@@
$\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\theta_{*}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\text{out}}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi_{\text{out}}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\theta_{*}}^{\eta}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\theta_{*}}^{\eta}(a|x)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\text{out}}}\bigg{[}\log\frac{\pi_{0}(a|x)\cdot\exp\bigl{(}\eta R(\theta_{*},x,a)\bigr{)}}{\pi_{\text{out}}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]},$
where the last equality follows from the fact that $\pi_{\theta_{*}}^{\eta}\propto\pi_{0}(a|x)\cdot\exp(\eta R(\theta_{*},x,a))$. To bound the suboptimality gap, we further have
$\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$=\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{E}_{a\sim\pi_{\text{out}}(\cdot|x)}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}$
$\geq\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\frac{1}{M}\sum_{x\in\mathcal{X}}\mathbb{P}_{\theta}(\pi_{\text{out}}(\theta(x))\leq 1/2)\cdot\biggl{[}\frac{1}{2}\log\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\log\frac{1+\exp(\eta c)}{2}\biggr{]}$
$\geq\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/4M}\Bigr{)}\Bigl{[}\frac{1}{2}\log\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\log\frac{1+\exp(\eta c)}{2}\Bigr{]}$ (D.1)
@@
$\frac{\text{d}}{\text{d}u}\Bigl{[}\frac{1}{2}\log\frac{1+e^{-u}}{2}+\frac{1}{2}\log\frac{1+e^{u}}{2}\Bigr{]}\Big{|}_{u=0}$ $=\frac{1}{2}\Bigl{[}\frac{1}{1+\exp(-u)}-\frac{1}{1+\exp(u)}\Bigr{]}\Big{|}_{u=0}=0,$
$\frac{\text{d}^{2}}{\text{d}u^{2}}\Bigl{[}\frac{1}{2}\log\frac{1+e^{-u}}{2}+\frac{1}{2}\log\frac{1+e^{u}}{2}\Bigr{]}$ $=\frac{\exp(u)}{[1+\exp(u)]^{2}}.$
@@
$\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{2}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/4M}\Bigr{)}\eta^{2}c^{2}\cdot\frac{1}{3+\exp(\eta c)}$
@@
$\frac{1}{2}\log\frac{1+\exp(-\eta c)}{2}+\frac{1}{2}\log\frac{1+\exp(\eta c)}{2}$ $=\frac{1}{2}\log\frac{2+\exp(\eta c)+\exp(-\eta c)}{4}$
$\geq\frac{1}{2}\cdot\frac{1}{2}\Bigl{(}\log\frac{\exp(\eta c)+\exp(-\eta c)}{2}\Bigr{)}$
$\geq\frac{1}{4}(\eta c-\log 2),$ (D.2)
where the first inequality follows from Jensen’s inequality. Substituting (D.2) into (D.1), we have
$\epsilon\geq\frac{1}{\eta}\mathbb{E}_{\theta\sim\text{Unif}(\Theta)}\mathbb{E}_{\pi_{\text{out}}}\Bigl{[}\log\frac{\pi_{\text{out}}(a|x)}{\pi^{*}(a|x)}\Bigr{]}\geq\frac{1}{4}\cdot\Bigl{(}\frac{1}{2}-\sqrt{c^{2}T/4M}\Bigr{)}(\eta c-\log 2)\cdot\frac{1}{\eta}.$
@@
$\leq\log(1/\delta)+\sum_{i=1}^{s}\bigg{(}\sqrt{\sigma(R(\theta_{*},x_{i},a_{i}^{2})-R(\theta_{*},x_{i},a_{i}^{1}))\cdot\sigma(R(\theta,x_{i},a_{i}^{2})-R(\theta,x_{i},a_{i}^{1}))}$
$\quad\quad+\sqrt{\sigma(R(\theta_{*},x_{i},a_{i}^{1})-R(\theta_{*},x_{i},a_{i}^{2}))\cdot\sigma(R(\theta,x_{i},a_{i}^{1})-R(\theta,x_{i},a_{i}^{2}))}-1\bigg{)}$
@@
$\quad-\frac{1}{2}\sum_{i=1}^{s}\bigg{(}\sqrt{\sigma(R(\theta_{*},x_{i},a_{i}^{1})-R(\theta_{*},x_{i},a_{i}^{2}))}-\sqrt{\sigma(R(\theta,x_{i},a_{i}^{1})-R(\theta,x_{i},a_{i}^{2}))}\bigg{)}^{2}$
@@
where the second inequality holds due to $\log(1+r)\leq r$ for $r>-1$, the equality follows from the fact that $\sigma(r)+\sigma(-r)=1$ and the last inequality holds since $\sigma^{\prime}(r)=\sigma(r)\cdot(1-\sigma(r))\geq e^{-B}$ for all $r\in[-B,B]$. ∎
@@
By Lemmas D.1 and D.2, we have with probability at least $1-\delta$, for any $\theta\in\Theta$,
@@
00 0 $\leq\log(N_{\mathcal{R}}(\epsilon_{c})/\delta)-\frac{1}{8}e^{-B}\sum_{i=1}^{n}\big{(}[R(\theta,x_{i},a_{i}^{2})-R(\theta,x_{i},a_{i}^{1})]-[R(\theta_{*},x_{i},a_{i}^{2})-R(\theta_{*},x_{i},a_{i}^{1})]\big{)}^{2}$
$\quad+O(n\epsilon_{c}).$
@@
for all $\theta^{c}\in\Theta^{c}$. By Lemma E.2 and the definition of $\Theta^{c}$, we further have
@@
with probability at least $1-2\delta$ by Lemma D.3 and the union bound.
@@
for all $\theta^{c}\in\Theta^{c}$. By Lemma E.2 and the definition of $\Theta^{c}$, we further have
@@
$\leq O\bigg{(}\frac{1}{m}B^{2}\log(N_{\mathcal{R}}(\epsilon_{c})/\delta\bigg{)}$
$\quad+\frac{1}{m}\sum_{i=1}^{n}\big{(}[R(\widehat{\theta},\widetilde{x}_{i},\widetilde{a}_{i}^{2})-R(\widehat{\theta},\widetilde{x}_{i},\widetilde{a}_{i}^{1})]-[R(\theta_{*},\widetilde{x}_{i},\widetilde{a}_{i}^{2})-R(\theta_{*},\widetilde{x}_{i},\widetilde{a}_{i}^{1})]\big{)}^{2}+B\epsilon_{c}).$ (D.5)
@@
By Lemma D.4, we have with probability at least $1-2\delta$, there exists a mapping $b_{1}:\mathcal{X}\to\mathbb{R}$ such that